1. IN A TRIANGLE ABC EQUATION OF SIDES AB & AC ARE 2x +3y=29 &


x + 2y=16.


if the midpoint of BC is (5,6) find equation of BC ?


2.AD, BE CF ARE THE ALTITUDES OF TRI. ABC WITH A(-4,5) D(3.2,


-4.6) E(4,1) F(-1,-4).


FIND EQUATION OF SIDES OF TRI. ABC & COORDINATES OF B & C?


SIR PLZ REPLY.. 


Grade:11

1 Answers

Rathod Shankar AskiitiansExpert-IITB
13 years ago

1)  say  C=(a,b)

C has to satisfy AC eqn

use this for AC equation    we get        a+2b=16.....................1

and we know BC's  mid point is (5,6)

say D=(5,6)

 

now use the equation of line which passes through two points. here our two points are C(a,b)  and D(5,6)

line is  

y - y1 = [(y2 - y1) / (x2 - x1)]*(x - x1)

y - 6 = [(b - 6) / (a - 5)]  * (x - 5)

 

this line has to pass throghu B(c,d) we get

c - 6 = [(b - 6) / (a - 5)]  * (d - 5)...........................2

 

 

again use it for B(c,d) and D(5,6)

using AB equation u will get    2c+3d=32.............3

again use the equation of line which passes through two points. here our two points are B(c,d)  and D(5,6)

line is  

y - y1 = [(y2 - y1) / (x2 - x1)]*(x - x1)

y - 6 = [(d - 6) / (c - 5)]  * (x - 5)

this line has to pass throghu C(a,b) we get

a - 6 = [(d - 6) / (c - 5)]  * (b - 5)...........................4

 

 now 4 equations 4 unknowns solve them and get the points B and D..after use two point line formula again we get the equations

 

2)

use same line rules and equation mentioned above u will get this answer too. do practice and solve second one so that u will remember it

 

Now you can win exciting gifts by answering the questions on
Discussion Forum. So help discuss any query on askiitians forum and
become an Elite Expert League askiitian.
Now you score 5+15 POINTS by uploading your Pic and Downloading the
Askiitians Toolbar : Click here to download the toolbar..